# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ 2. A priori estimates and technical results
Let $\rho_{\varepsilon}$ be as stated, then
$\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(t)=e^{-\varepsilon t}\int_{\mathbb{R}^{d}}\rho_{0}.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)=$ $\;(1+\varepsilon)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\Delta\varphi\,{\rm d}x\,{\rm d}t+\chi\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\frac{\nabla\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\nabla\varphi\,{\rm d}x\,{\rm d}t$ (12)
$\,-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{-}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x.$
But notice that
-- $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{-}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t=\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$\leq\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\chi\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\rho_{\varepsilon}\frac{\nabla\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$\leq{(1+\chi)}\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\rho_{\varepsilon}|\nabla\varphi|\,{\rm d}x\,{\rm d}t.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq$ $(1+\varepsilon)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\Delta\varphi\,{\rm d}x\,{\rm d}t+(2\chi+1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\varphi\right|\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x$
$=$ $\;(1+\varepsilon)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\frac{\alpha(\frac{\alpha}{2}+1)\frac{|x|^{2}}{R^{4}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+2}}-\frac{\alpha d\frac{1}{R^{2}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\right]\,{\rm d}x\,{\rm d}t$
$+3\chi\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}\alpha\frac{|x|}{R^{2}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x$
$\leq$ $\;(1+\varepsilon)\frac{\alpha}{2R}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}(\frac{\alpha}{2}+1-d)}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\,{\rm d}x\,{\rm d}t+3\chi\frac{\alpha}{R}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}}}$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq\;\frac{C}{R}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x,$ (13)
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq e^{\frac{CT}{R}}\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x\leq e^{\frac{CT}{R}}\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x.$
By monotone convergence we can take the limit $R\rightarrow\infty$ and conclude that $\int\rho_{\varepsilon}\leq\int\rho_{0}$.
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x)(T)=-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x+\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x.$ (14)
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x)(T)=e^{-\varepsilon T}\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x,$
@@
Proposition 1 ($L^{p}$ estimates).
Let $\rho_{\varepsilon}$ be as stated, then
$\|\rho_{\varepsilon}\|_{L^{\infty}((0,T);L^{p}(\mathbb{R}^{d}))}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})},\qquad\|\nabla\rho_{\varepsilon}^{\frac{p}{2}}\|_{L^{2}((0,T)\times\mathbb{R}^{d})}\leq C,$
where $C$ only depends on $\|\rho_{0}\|_{L^{p}},p$ and $\chi$.
@@
First, notice that testing (11) with $\rho_{\varepsilon}^{p-1}$we get that $\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}(t)\leq\int_{\mathbb{R}^{d}}\rho_{0}^{p}$. Then, we compute
$\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}(t)\leq$ $-(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-2}\lvert\nabla\rho_{\varepsilon}\rvert^{2}+\chi(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-1}\lvert\nabla\rho_{\varepsilon}\rvert+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{0}^{p}$
$\leq$ $-\frac{(p-1)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-2}\lvert\nabla\rho_{\varepsilon}\rvert^{2}+\chi^{2}(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{0}^{p}$
$\leq$ $-\frac{2(p-1)}{p^{2}}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla\rho_{\varepsilon}^{\frac{p}{2}}\rvert^{2}+\chi^{2}(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{0}^{p}+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{0}^{p}.$
We get that
$\|\rho_{\varepsilon}\|^{p}_{L^{\infty}(0,T;L^{p}(\mathbb{R}^{d}))}+\|\nabla\rho_{\varepsilon}^{\frac{p}{2}}\|_{L^{2}(0,T;L^{2}(\mathbb{R}^{d}))}\leq C,$
@@
Let $\rho_{\varepsilon}$ be as stated, then
$\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(t)(1+|x|^{2})\,{\rm d}x\leq e^{C\;t}\int_{\mathbb{R}^{d}}\rho_{0}(1+|x|^{2})\,{\rm d}x,$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x)(T)\leq C\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}(1+|x|^{2})\varphi\,{\rm d}x.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x)(T)\leq e^{C\;t}\int_{\mathbb{R}^{d}}\rho_{0}(1+|x|^{2})\varphi\,{\rm d}x.$
@@
Let $\rho_{\varepsilon}$ be as stated, then
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x)(T)+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\lvert\nabla\rho_{\varepsilon}\rvert^{2}}{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t\leq\bar{C}[\rho_{0}],$
@@
Let us test (11) with the test function $\varphi=\log(\rho_{\varepsilon}+\sigma)-\log{\sigma}$, with $10<\sigma<<$ 1. After some computations we get
$\left(\int_{\mathbb{R}^{d}}\left[(\rho_{\varepsilon}+\sigma)\log(\rho_{\varepsilon}+\sigma)-\rho_{\varepsilon}\log\sigma\right]\,{\rm d}x\right)(T)+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\log(\rho_{\varepsilon}+\sigma)-\log{\sigma}\right]\,{\rm d}x\,{\rm d}t$
$\qquad\leq\int_{\mathbb{R}^{d}}\left[(\rho_{0}+\sigma)\log(\rho_{0}+\sigma)-\rho_{0}\log\sigma\right]\,{\rm d}x,$
@@
$\left(\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\log(\rho_{\varepsilon}+\sigma)\,{\rm d}x\right)(T)\leq$ $\int_{\mathbb{R}^{d}}(\rho_{0}+\sigma)\log(\rho_{0}+\sigma)\,{\rm d}x$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log{(\rho_{\varepsilon}+\sigma)}\,{\rm d}x\,{\rm d}t$
$+\log\sigma\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x-\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t\right).$
@@
$\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x\right)(T)\leq$ $\int_{\mathbb{R}^{d}}\rho_{0}\log\rho_{0}\,{\rm d}x-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t.$
@@
$\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}|\log\rho_{\varepsilon}|\,{\rm d}x\right)(T)\leq e^{\varepsilon T}C[\rho_{0}],$
where $C[\rho_{0}]$ is a constant that depends on $\int_{\mathbb{R}^{d}}\rho_{0}|\log{\rho_{0}}|\,{\rm d}x$, the mass and the second moment of $\rho_{0}$.
@@
$h[\rho_{\varepsilon}](T)+\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\,{\rm d}x\,{\rm d}t\leq\frac{\chi^{2}}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}h[\rho_{\varepsilon}](t)\,{\rm d}t+h[\rho_{0}],$
@@
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\,{\rm d}x\,{\rm d}t\leq\frac{\chi^{2}}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t+e^{\varepsilon T}C[\rho_{0}].$
@@
Let $\rho_{\varepsilon}$ be as stated, then
$\|\nabla\rho_{\varepsilon}\|_{L^{p}((t,T)\times\mathbb{R}^{d})}<\frac{1}{t^{1/2}}C(\|\rho_{0}\|_{L^{p}}),\quad\forall\;t>0.$
@@
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}=\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right),$
@@
$g_{t}=(1+\varepsilon)\Delta g-\varepsilon g,\quad g(x,0)=\rho_{0}.$
@@
Let $\rho_{\varepsilon}$ be as stated, then for every $\tau>0$ it holds
$\|\rho_{\varepsilon}\|_{L^{\infty}({\tau},T;\mathbb{R}^{d})}\leq C{T^{\frac{1}{p}}\left(1+\tau^{-\frac{d+2}{2p}}\right)}\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})},$

### v3

@@ 2. A priori estimates
Let $\rho_{\varepsilon}$ be a weak solution to (12) for any $\varepsilon\geq 0$. It holds:
$\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(t)=e^{-\varepsilon t}\int_{\mathbb{R}^{d}}\rho_{in}.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)=$ $\;(1+\varepsilon)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\Delta\varphi\,{\rm d}x\,{\rm d}t+\chi\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\frac{\nabla\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\nabla\varphi\,{\rm d}x\,{\rm d}t$ (13)
$\,-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{-}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{in}\varphi\,{\rm d}x.$
Here we have rewritten
$\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}=\nabla\rho_{\varepsilon}-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\nabla\rho_{\varepsilon}-\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{-}\nabla\rho_{\varepsilon}.$ (14)
Notice that
- $\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{-}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t=\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$\leq\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\chi\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\rho_{\varepsilon}\frac{\nabla\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$\leq{(1+\chi)}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\rho_{\varepsilon}|\nabla\varphi|\,{\rm d}x\,{\rm d}t.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq$ $(1+\varepsilon)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\Delta\varphi\,{\rm d}x\,{\rm d}t+(2\chi+1)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\varphi\right|\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{in}\varphi\,{\rm d}x$
$=$ $\;(1+\varepsilon)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\frac{\alpha(\frac{\alpha}{2}+1)\frac{|x|^{2}}{R^{4}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+2}}-\frac{\alpha d\frac{1}{R^{2}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\right]\,{\rm d}x\,{\rm d}t$
$+3\chi\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}\alpha\frac{|x|}{R^{2}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{in}\varphi\,{\rm d}x$
$\leq$ $\;(1+\varepsilon)\frac{\alpha}{2R}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}(\frac{\alpha}{2}+1-d)}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\,{\rm d}x\,{\rm d}t+3\chi\frac{\alpha}{R}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}}}$
$-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{in}\varphi\,{\rm d}x.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq\;\frac{C}{R}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{in}\varphi\,{\rm d}x,$ (15)
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq e^{\frac{CT}{R}}\int_{\mathbb{R}^{d}}\rho_{in}\varphi\,{\rm d}x\leq e^{\frac{CT}{R}}\int_{\mathbb{R}^{d}}\rho_{in}\,{\rm d}x.$
By monotone convergence we can take the limit $R\rightarrow\infty$ and conclude that $\int\rho_{\varepsilon}\leq\int\rho_{in}$.
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x)(T)=-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x+\int_{\mathbb{R}^{d}}\rho_{in}\,{\rm d}x.$ (16)
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x)(T)=e^{-\varepsilon T}\int_{\mathbb{R}^{d}}\rho_{in}\,{\rm d}x,$
@@
Proposition 1 ($L^{p}$, $H^{1}$ estimates).
Let $\rho_{\varepsilon}$ be a weak solution to (12) for any $\varepsilon\geq 0$, then
$\|\rho_{\varepsilon}\|^{p}_{L^{\infty}((0,T);L^{p}(\mathbb{R}^{d}))}+\|\nabla\rho_{\varepsilon}^{\frac{p}{2}}\|^{2}_{L^{2}((0,T)\times\mathbb{R}^{d})}\leq C,$
where $C$ only depends on $\|\rho_{in}\|_{L^{p}},p$, $T$, and $\chi$.
@@
First, notice that testing (12) with $\rho_{\varepsilon}^{p-1}$we get that $\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}(t)\leq\int_{\mathbb{R}^{d}}\rho_{in}^{p}$. Then, we compute
$\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}(t)\leq$ $-(p-1)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-2}\lvert\nabla\rho_{\varepsilon}\rvert^{2}+\chi(p-1)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-1}\lvert\nabla\rho_{\varepsilon}\rvert+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{in}^{p}$
$=$ $-\frac{2(p-1)}{p^{2}}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\lvert\nabla\rho_{\varepsilon}^{\frac{p}{2}}\rvert^{2}+\chi^{2}(p-1)\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{in}^{p},$
which yields
$\|\rho_{\varepsilon}\|^{p}_{L^{\infty}(0,T;L^{p}(\mathbb{R}^{d}))}+\|\nabla\rho_{\varepsilon}^{\frac{p}{2}}\|^{2}_{L^{2}(0,T;L^{2}(\mathbb{R}^{d}))}\leq C,$
@@
Let $\rho_{\varepsilon}$ be a weak solution to (12) for any $\varepsilon\geq 0$, then
$\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(t)(1+|x|^{2})\,{\rm d}x\leq e^{C\;t}\int_{\mathbb{R}^{d}}\rho_{in}(1+|x|^{2})\,{\rm d}x,$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x)(T)\leq C\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{in}(1+|x|^{2})\varphi\,{\rm d}x.$
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x)(T)\leq e^{CT}\int_{\mathbb{R}^{d}}\rho_{in}(1+|x|^{2})\varphi\,{\rm d}x.$
@@
Let $\rho_{\varepsilon}$ be a weak solution to (12) for any $\varepsilon\geq 0$, then
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x)(T)+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\lvert\nabla\rho_{\varepsilon}\rvert^{2}}{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t\leq\bar{C}[\rho_{in}],$
@@
Let us test (12) with the test function $\varphi=\log(\rho_{\varepsilon}+\sigma)-\log{\sigma}$, with $0<\sigma<<1$. After some computations we get
$\left(\int_{\mathbb{R}^{d}}\left[(\rho_{\varepsilon}+\sigma)\log(\rho_{\varepsilon}+\sigma)-\rho_{\varepsilon}\log\sigma\right]\,{\rm d}x\right)(T)+\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\log(\rho_{\varepsilon}+\sigma)-\log{\sigma}\right]\,{\rm d}x\,{\rm d}t$
$\qquad\leq\int_{\mathbb{R}^{d}}\left[(\rho_{in}+\sigma)\log(\rho_{in}+\sigma)-\rho_{in}\log\sigma\right]\,{\rm d}x,$
@@
$\left(\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\log(\rho_{\varepsilon}+\sigma)\,{\rm d}x\right)(T)\leq$ $\int_{\mathbb{R}^{d}}(\rho_{in}+\sigma)\log(\rho_{in}+\sigma)\,{\rm d}x$
$-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log{(\rho_{\varepsilon}+\sigma)}\,{\rm d}x\,{\rm d}t$
$+\log\sigma\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x-\int_{\mathbb{R}^{d}}\rho_{in}\,{\rm d}x+\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t\right).$
@@
$\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x\right)(T)\leq$ $\int_{\mathbb{R}^{d}}\rho_{in}\log\rho_{in}\,{\rm d}x-\varepsilon\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t.$
@@
$\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}|\log\rho_{\varepsilon}|\,{\rm d}x\right)(T)\leq e^{\varepsilon T}C[\rho_{in}],$
where $C[\rho_{in}]$ is a constant that depends on $\int_{\mathbb{R}^{d}}\rho_{in}|\log{\rho_{in}}|\,{\rm d}x$, the mass and the second moment of $\rho_{in}$.
@@
$h[\rho_{\varepsilon}](T)+\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\,{\rm d}x\,{\rm d}t\leq\frac{\chi^{2}}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}h[\rho_{\varepsilon}](t)\,{\rm d}t+h[\rho_{in}],$
@@
$\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\,{\rm d}x\,{\rm d}t\leq\frac{\chi^{2}}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t+e^{\varepsilon T}C[\rho_{in}].$
@@
Let $\rho_{\varepsilon}$ be a weak solution to (12) for any $\varepsilon\geq 0$, then
$\|\nabla\rho_{\varepsilon}\|_{L^{p}((t,T)\times\mathbb{R}^{d})}<\frac{1}{t^{1/2}}C(\|\rho_{in}\|_{L^{p}}),\quad\forall\;t>0.$
@@
[…]
@@
$g_{t}=(1+\varepsilon)\Delta g-\varepsilon g,\quad g(x,0)=\rho_{in}.$
@@
Let $\rho_{\varepsilon}$ be a weak solution to (12) for any $\varepsilon\geq 0$, then for every $t>0$ it holds
$\|\rho_{\varepsilon}\|_{L^{\infty}({t},T;\mathbb{R}^{d})}\leq C{T^{\frac{1}{p}}\left(1+t^{-\frac{d+2}{2p}}\right)}\|\rho_{in}\|_{L^{p}(\mathbb{R}^{d})},$